car moves at a speed of 5m/s. It accelerates uniformly to a new speed of 15m/s. If this takes 30 seconds, calculate the distance that the car travels.


1
Expert's answer
2021-05-09T13:04:36-0400
"s=\frac{v_i+v_f}{2}t\\s=\frac{5+15}{2}30=300\ m"